OpenStudy (anonymous):

what is 3 9ths reduced? (the fraction)

OpenStudy (anonymous):

3/9=1/3

OpenStudy (help):

1/3

OpenStudy (anonymous):

ok thx but can u show the work plzz

OpenStudy (anonymous):

sure one sec and I'll write it out

OpenStudy (anonymous):

ok thx

OpenStudy (anonymous):

3/9 what can go into both 3 and 9? the answer is three so divide both 3 and 9 by 3 3/3=1 9/3=3 your answer is 1/3

OpenStudy (anonymous):

I is 1/3

OpenStudy (anonymous):

I mean that it is 1/3

OpenStudy (anonymous):

ok thx

OpenStudy (anonymous):

no problem anytime:)

OpenStudy (anonymous):

oh wait but u have to divde dont u?

OpenStudy (anonymous):

not multiply?

OpenStudy (anonymous):

yes that's what I did...

OpenStudy (anonymous):

oh ok srry

OpenStudy (anonymous):

That's ok do you understand it?